# Kartik Narayan, Vibashan VS, Vishal M. Patel

### v1

@@ B.3 Source Datasets
B.4 Dataset Statistics
The FaceXBench dataset is derived from 25 public datasets and one newly created dataset. The number of questions sourced from each dataset, along with the type of questions (multiple images, single images, or text-only), as well as the associated tasks and categories, are summarized in Table B.1.
@@ B.4 Dataset Statistics
Table B.1: Question distribution of FaceXBench across datasets
B.5 Images used in the dataset
Figure B.1 displays a subset of the facial images used in the dataset, highlighting the diversity of faces included in the benchmark. The benchmark consists of face images with varying backgrounds, resolutions, and head pose orientations, as well as a wide range of facial expressions. It includes individuals from different age groups, genders, and racial backgrounds, with each face characterized by a versatile set of attributes.
Figure B.1: Collage of a subset of images from the dataset, showcasing the diversity of images used in FaceXBench.
Appendix C Dataset Collection
@@ C.1 Question Templates
C.2 Generating Distractor Options
@@
C.3 FaceXAPI dataset prompt
@@
C.4 Option Extraction from Prediction
@@ C.5 Dataset Format
C.6 Other Evaluation Settings
@@ D.1 Baselines
Table D.1: Vision encoder, LLM and LLM size of the baseline models
D.2 Results Across Tasks
To provide additional insights, we summarize the task-wise performance of various models in Table D.2. We observe that models struggle with tasks such as head pose estimation, crowd counting, and deepfake detection. Conversely, tasks like gender prediction, facial expression recognition, and high-resolution face recognition are comparatively easier. Additionally, we present the results of selected models under the ”in-context” and ”chain-of-thought” evaluation settings in Table D.4 and Table D.4, respectively. We observe a consistent drop in performance, indicating that the models are not capable of effectively processing in-context information and fail to reason in tasks related to face understanding.
@@ D.2 Results Across Tasks
Frequent Choice 32.22 30.00 29.40 31.00 34.67 28.67 28.75 35.00 46.00 40.00 36.50 31.80 28.75 27.00 32.33
@@
Table D.2: Performance of various models across all evaluation tasks of FaceXBench.
@@
Table D.3: (a) Performance of models in in-context evaluation setting.
@@
Table D.4: (b) Performance of models in chain-of-thought evaluation setting.
D.3 Failure Cases
@@
Appendix E Ethical Considerations

### v2

@@ C.3 Source Datasets
C.4 Dataset Statistics
The FaceXBench dataset is derived from 25 public datasets and one newly created dataset. The number of questions sourced from each dataset, along with the type of questions (multiple images, single images, or text-only), as well as the associated tasks and categories, are summarized in Table C.1.
@@ C.4 Dataset Statistics
Table C.1: Question distribution of FaceXBench across datasets
C.5 Images used in the dataset
Figure C.1 displays a subset of the facial images used in the dataset, highlighting the diversity of faces included in the benchmark. The benchmark consists of face images with varying backgrounds, resolutions, and head pose orientations, as well as a wide range of facial expressions. It includes individuals from different age groups, genders, and racial backgrounds, with each face characterized by a versatile set of attributes.
Figure C.1: Collage of a subset of images from the dataset, showcasing the diversity of images used in FaceXBench.
Appendix D Dataset Collection
@@ D.1 Question Templates
D.2 Generating Distractor Options
@@
D.3 FaceXAPI dataset prompt
@@
D.4 Option Extraction from Prediction
@@ D.5 Dataset Format
D.6 Other Evaluation Settings
@@ E.1 Baselines
Table E.1: Vision encoder, LLM and LLM size of the baseline models
E.2 Results Across Tasks
To provide additional insights, we summarize the task-wise performance of various models in Table E.2. We observe that models struggle with tasks such as head pose estimation, crowd counting, and deepfake detection. Conversely, tasks like gender prediction, facial expression recognition, and high-resolution face recognition are comparatively easier. Additionally, we present the results of selected models under the ”in-context” and ”chain-of-thought” evaluation settings in Table E.4 and Table E.4, respectively. We observe a consistent drop in performance, indicating that the models are not capable of effectively processing in-context information and fail to reason in tasks related to face understanding.
@@ E.2 Results Across Tasks
Frequent Choice 26.68 27.50 27.40 27.20 28.00 28.67 26.00 35.00 41.25 40.00 33.50 28.20 26.25 26.50 28.00
Human 70.28 77.75 47.00 75.20 40.33 32.67 91.75 59.00 71.75 94.00 74.50 94.80 60.75 94.00 58.33
Vision SOTA models 84.50 91.00 68.00 88.60 59.33 63.00 98.00 74.00 90.25 57.00 91.75 96.40 81.25 96.75 88.67
@@
Table E.2: Performance of various models across all evaluation tasks of FaceXBench.
@@
Table E.3: (a) Performance of models in in-context evaluation setting.
@@
Table E.4: (b) Performance of models in chain-of-thought evaluation setting.
E.3 Human and SOTA Vision models evaluation details
For human evaluation, we conducted assessments with five graduate student participants and reported the average results. For evaluating Vision SOTA models, certain tasks, such as tool retrieval and celebrity identification, do not have a specific benchmark. In these cases, we considered Gemini Pro 1.5 [99] as the SOTA model. For tasks such as expression recognition, age, gender, and race estimation, attribute prediction, and head pose estimation, we use FaceXFormer [75] as the SOTA model. For high-resolution face recognition, we employ ArcFace [20], while for low-resolution face recognition, we use PETALface [74]. Additionally, we utilize SegFace [76] for face parsing, FLIP [43] for face anti-spoofing, and XceptionNet [17] for deepfake detection.
As can be seen from Table E.2, there are several tasks where human participants perform significantly worse than SOTA models, primarily due to the nature of the tasks. These include deepfake detection, head pose estimation, celebrity identification, crowd counting, and age estimation. In deepfake detection, some manipulations are more pronounced in the spectral domain but remain imperceptible to the human eye. Similarly, age estimation requires high precision, as the distractor options are designed to be very close in value, meaning even a small human error results in an incorrect choice, making the benchmark more challenging and forcing the model to be precise. For head pose estimation, humans struggle to visualize fine angular changes in yaw, pitch, and roll, leading to poor performance. Additionally, the low human performance in celebrity identification is primarily due to the participants’ limited familiarity with a wide range of celebrities.
E.4 Failure Cases
@@
Appendix F Ethical Considerations